OpenStudy (anonymous):

rewrite 6log5x^3 in a from that does not use exponents

OpenStudy (anonymous):

6log(5)+6log(x)

OpenStudy (anonymous):

this is the best........i feel

OpenStudy (mertsj):

\[\log_{10}5x^3=\log_{10}5+\log_{10}x^3=\log_{10}5+3\log_{10}x \] \[6\log_{10}5x^3=6(\log_{10}5+3\log_{10}x)= 6\log_{10}5+18\log_{10}x \]